Find the polar form of complex number (2+2isqrt(3))
1
Expert's answer
2019-10-16T11:23:32-0400

"z=2+2\sqrt{3}i."

"r=\sqrt{2^2+(2\sqrt{3})^2}=\sqrt{16}=4."

"\phi=arctan(\frac{2\sqrt{3}}{2})=arctan(\sqrt{3})=\frac{\pi}{3}."

Polar form: "z=4(cos(\frac{\pi}{3})+isin(\frac{\pi}{3}))".